OpenStudy (anonymous):

Is -5x-2 the same as 2-5x?

OpenStudy (anonymous):

you have -5x and -2 ... so -5x-2 = -2-5x

OpenStudy (anonymous):

Then how would you rewrite 2-5x? Just 5x-2?

OpenStudy (anonymous):

you have +2 and -5x this time,so , +2-5x = -5x+2

OpenStudy (anonymous):

Oh! Okay, yeah I had a dumb moment lol. Thank you!

OpenStudy (anonymous):

Welcome :)